On July 1 of the current year F purchased a franchise and commenced operating a new business as a sole proprietor. The franchise cost $100,000 and has a 10-year term. What is the maximum CCA deduction that F can claim in the current year?
Answer #1

Capital cost allowance ( CCA Deduction ) -

Franchise Cost = $ 100000

term - 10 years

Current year claim = total cost / no of year = 100000/10 = $10000

remaining period in current year = 6 month ( july to december )

So current year maximum CCA deduction = 10000/12* 6 = $ 5000